Forty eight grams of carbon is added and reacted to 134grams of oxygen to produce Carbon dioxide. Find (a) the balance chemical equation (4points), (b) the mass of the product (4points), (c) the limiting reactant (4points), (d) the mass of excess reactant not used in the reaction (6points). Write the complete formula and solution for each problem.You can use a separate sheet if needed. Use whole numbers only.

1
Expert's answer
2022-03-25T06:49:02-0400

a) C + O2 = CO2

b) Mass = moles * molecular mass

Mole ratio between C:CO2 = 1:1

C moles=4 moles

Mass of CO2= 44 * 4 = 176 grams

c)Limiting reagent is carbon because it has less moles

d) mass of excess reactant = 0.1875 * 32 = 6 grams